OpenStudy (anonymous):

One-fourths of a herd of camels was seen in the forest.Twice the square root of the number of herd had gone to mountains and the remaining 15 camels were seen on the bank of a river.find the total number of camels.

OpenStudy (anonymous):

Putting all these into an equation we can work with: (let the total no. of camels be x) \[\frac{ x }{ 4 } + 2 \sqrt x + 15 = x\]Collecting like terms...\[\frac{ 3 }{ 4 }x + 2 \sqrt x + 15 = 0\]

OpenStudy (anonymous):

\[x = 36\]

OpenStudy (anonymous):

That's the total no. of camels.

OpenStudy (anonymous):

Welcome to OS! :)

OpenStudy (anonymous):

thank u soooo much

OpenStudy (anonymous):

x =36

OpenStudy (anonymous):

Here's how you solve for x:\[-\frac{ 3 }{ 4 }x+2 \sqrt x + 15 = 0\]LCD = 4, multiply through to remove fraction\[-3x + 8 \sqrt x + 60 = 0\]\[8 \sqrt x = 3x - 60\]Square both sides to remove radical\[9x^2 -424x + 3600\]Factorize\[9x^2 -100x-324x +3600=0 \rightarrow x(9x-100)-36(9x-100)=0\]\[(x-36)(9x-100)=0\]Using the zero factor principle:\[x-36 =0 , 9x -100 = 0 \rightarrow x = 36, \frac{ 100 }{ 9 }\]So, we see that there are two values of x, but we need just one. Which do we choose? The one which makes more sense within the context of our problem.\[x = 36\]I hope this is clear enough.

OpenStudy (anonymous):

yes.it's more than enough.thank you soo much.